2/3 - 6x = 10 what is x?
Answer by tracyc0912(40)

2/3 - 6x=10 Multiply whole thing by 3 to remove your fraction
2-18x=30 Subtract 2 from each side
-18x= 28 Divide both sides by -18
x= -28/18 Reduce your fraction
x=-14/9 Double check by putting this figure back into the original to make sure it works out correctly!
Good Luck :-)